The probability that a freshman entering AAU (Science Faculty) will survive first semester is 0.92. (From


1993/94 academic year statistics). Assuming this pattern remain unchanged over the subsequent years, what


is the probability that among 100 randomly selected freshmen in first semester,


a) None will survive?


b) Exactly 97 will survive?


c) At least three will survive?

1
Expert's answer
2022-04-19T02:57:52-0400

"p=0.92\\X\sim Bin\left( 100,0.92 \right) \\a:\\P\left( X=0 \right) =\left( 1-0.92 \right) ^{100}=2.03704\times 10^{-110}\\b:\\P\left( X=97 \right) =C_{100}^{97}\cdot 0.92^{97}\cdot \left( 1-0.92 \right) ^3=0.0254331\\c:\\P\left( X\geqslant 3 \right) =1-P\left( X=0 \right) -P\left( X=1 \right) -P\left( X=2 \right) =\\=1-2.03704\times 10^{-110}-C_{100}^{1}\cdot 0.92^1\cdot \left( 1-0.92 \right) ^{99}-C_{100}^{99}\cdot 0.92^2\cdot \left( 1-0.92 \right) ^{98}=\\=1-2.92804\times 10^{-106}"